Klean Fiber Company is the creator of Y-Go, a technology that weaves silver into its fabrics to kill bacteria and odor on clothing while managing heat. Y-Go has become very popular in undergarments for sports activities. Operating at capacity, the company can produce 1,087,000 Y-Go undergarments a year. The per unit and the total costs for an individual garment when the company operates at full capacity are as follows. Per Undergarment Total Direct materials $1.96 $2,130,520 Direct labor 0.41 445,670 Variable manufacturing overhead 0.99 1,076,130 Fixed manufacturing overhead 1.58 1,717,460 Variable selling expenses 0.30 326,100 Totals $5.24 $5,695,880 The U.S. Army has approached Klean Fiber and expressed an interest in purchasing 249,700 Y-Go undergarments for soldiers in extremely warm climates. The Army would pay the unit cost for direct materials, direct labor, and variable manufacturing overhead costs. In addition, the Army has agreed to pay an additional $1.09 per undergarment to cover all other costs and provide a profit. Presently, Klean Fiber is operating at 70% capacity and does not have any other potential buyers for Y-Go. If Klean Fiber accepts the Army’s offer, it will not incur any variable selling expenses related to this order. Prepare an incremental analysis for the Klean Fiber. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Answer #1
Reject Order Accept Order Net Income Increase
(Decrease)
Revenues 0 1111165 1111165
Variable costs:
Direct materials 0 -489412 -489412
Direct labor 0 -102377 -102377
Variable overhead 0 -247203 -247203
Total variable costs 0 -838992 -838992
Net income 0 272173 272173
Workings:
Revenues =249700*(1.96+0.41+0.99+1.09)
Variable costs:
Direct materials =249700*1.96
Direct labor =249700*0.41
Variable overhead =249700*0.99